OpenStudy (anonymous):

Hello I need help solving this: find all points having an x-coordinate of 1 whose distance from the point (-2 -3) is 5... any suggestions??

OpenStudy (anonymous):

(1,y) and (-2,-3) Distance formula. \[d^2=(1-(-2))^2+(y-(-3))^2=3^2+(y+3)^2\] \[d^2=9+y^2+6y+9\] \[d=5\] \[25=y^2+6y+18\] \[y^2+6y-7=0\] \[(y+7)(y-1)=0\] y is either -7 or 1.

OpenStudy (anonymous):

That's what I got too but I wasn't sure :-/ Thank you so much! I really appreciate it!